\begin{thebibliography}{999}

\bibitem{Nisar-Saiful} K S Nisar, S R Mondal and J Choi, Certain inequalities involving the $\mathtt{k}$-Struve function, submitted, 2016

\bibitem{ABRAMOWITZ} M. Abramowitz and I. A. Stegun, A Handbook of Mathematical Functions, New York, (1965).
\bibitem{Andrews-Askey} G. E. Andrews, R. Askey and R. Roy, Special functions, Encyclopedia of Mathematics
and its Applications, 71, Cambridge Univ. Press, Cambridge, 1999.
\bibitem{Romero-Cerutti} L.G. Romero, G.A.Dorrego and R.A. Cerutti, The k-Bessel function of first kind, International Mathematical forum, 38(7)(2012), 1859--1854.

\bibitem{Watson} GN. Watson, A Treatise on the Theory of Bessel Functions, Cambridge Mathematical Library Edition, Camdridge University Press, Camdridge (1995). Reprinted (1996)

\bibitem{Erdélyi-1} A. Erd\'{e}lyi, W. Magnus, F. Oberhettinger\ and\ F.G. Tricomi, Higher transcendental functions, I, II, McGraw-Hill
Book Company, Inc., New York, 1953. NewYork, Toronto, London, 1953.

\bibitem{Diaz} R. Diaz and E. Pariguan, On hypergeometric functions and k-Pochhammer symbol, Divulgaciones Matematicas 15(2) (2007), 179--192

\bibitem{Kwara14} Kwara Nantomah, Edward Prempeh, Some Inequalities for the k-Digamma Function, Mathematica Aeterna, 4(5) (2014), 521--525.

\bibitem{Mubeen13} S. Mubeen, M. Naz and G. Rahman, A note on k-hypergemetric differential equations, Journal of Inequalities and Special
Functions, 4(3) (2013), 8--43.

\bibitem{Biernacki-Krzy} M. Biernacki and J. Krzy\.z, On the monotonity of certain functionals in the theory of analytic functions,
Ann. Univ. Mariae Curie-Sk lodowska. Sect. A. 9  (1957), 135--147.

\bibitem{Kilbas-frac} A. A. Kilbas\ and\ N. Sebastian, Fractional
integration of the product of Bessel function of the first kind, Fract.
Calc. Appl. Anal. \textbf{13} (2010), no.~2, 159--175.


\bibitem{CGK} C. G. Kokologiannaki, Properties and inequalities of generalized $k$-gamma, beta and zeta functions, Int. J. Contemp. Math. Sci.  5(13-16) (2010), 653--660.

\bibitem{CGK2} C. G. Kokologiannaki\ and\ V. Krasniqi, Some properties of the $k$-gamma function, Matematiche (Catania), 68(1) (2013), 13--22.

\bibitem{VK} V. Krasniqi, A limit for the $k$-gamma and $k$-beta function, Int. Math. Forum,  5(33-36) (2010), 1613--1617.

\bibitem{MM}    M. Mansour, Determining the $k$-generalized gamma function $\Gamma\sb k(x)$ by functional equations, Int. J. Contemp. Math. Sci., 4(21-24) (2009), 1037--1042.

\bibitem{Askey} G. E. Andrews, R. Askey\ and\ R. Roy,  Special functions, Cambridge Univ. Press, Cambridge, 1999.

\bibitem{CFox} C. Fox, The asymptotic expansion of generalized hypergeometric functions, Proc. London. Math. Soc.,  27(4) (1928), 389--400.

\bibitem{Kilbas} A. A. Kilbas, M. Saigo\ and\ J. J. Trujillo, On the generalized Wright function, Fract. Calc. Appl. Anal.  5(4) (2002), 437--460.

\bibitem{Kilbas-itsf} A. A. Kilbas\ and\ N. Sebastian, Generalized fractional integration of Bessel function of the first kind,
Integral Transforms Spec. Funct.  19 (11-12) (2008), 869--883.

\bibitem{KST} A. A. Kilbas, H. M. Srivastava, and J. J. Trujillo, Theory and Applications of Fractional
Differential Equations, North-Holland Mathematics Studies 204, Elsevier, Amsterdam, 2006.

\bibitem{Rainville} E. D. Rainville,  Special functions, Macmillan, New York, 1960.

\bibitem{Wright-2} E. M. Wright, The asymptotic expansion of integral functions defined by Taylor series,
Philos. Trans. Roy. Soc. London, Ser. A. 238 (1940), 423--451.

\bibitem{Wright-3} E. M. Wright, The asymptotic expansion of the generalized hypergeometric function, Proc. London Math. Soc. (2) 46(1940), 389--408.
\bibitem{Mi} D.S. Mitrinovi\'c, {\em Analytic inequalities}, Springer, New York, 1970.

\bibitem{Nair-1} S. S. Nair: Pathway fractional integration operator, Fract.
Calc. Appl. Anal., 12(3) (2009), 237-252.

\bibitem{Mathai-pathway} A. M. Mathai : A pathway to matrix-variate gamma
and normal densities, Linear Algebra Appl., 396 (2005), 317-328.

\bibitem{Mathai-Habold-1} A. M. Mathai, H. J. Haubold: On generalized
distributions and path-ways, Phys. Lett. A, 372 (2008), 2109-2113.

\bibitem{Mathai-Habold-2} A. M. Mathai, H. J. Haubold: Pathway model,
superstatistics, Tsallis statistics and a generalize measure of entropy,
Phys. A, 375 (2007), 110-122.

\bibitem{Praveen-pathway} Dumitru Baleanu, Praveen Agarwal: A Composition
Formula of the Pathway Integral Transform Operator,Note di Matematica,Note
Mat. 34 (2014) no. 2, 145--155.

\bibitem{Purohit} P. Agarwal and S.D. Purohit,The Unified Pathway Fractional
Integral Formulae, Fract. Calc. Appl., 4(9) (2013), 1-8.

\end{thebibliography}

\documentclass[a4paper,12pt,reqno]{amsart}
\usepackage{amsmath,amsthm,amssymb}
\usepackage{graphicx, color}
\usepackage[numbers,sort&compress]{natbib}
\nonstopmode \numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]

 \newtheorem{corollary}{Corollary}[section]

\newtheorem{lemma}{Lemma}[section]

\allowdisplaybreaks

\allowdisplaybreaks

\begin{document}
\title[Pathway of $\mathtt{k}$-Struve function]{ Pathway fractional integral operators involving  $\mathtt{k}$-Struve function}

\author{Kottakkaran.S.Nisar}
\address{Department of Mathematics, College of Arts and Science-Wadi Al dawser,11991,\\
Prince Sattam bin Abdulaziz University, Saudi Arabia}
\email{ksnisar1@gmail.com, n.sooppy@psau.edu.sa}

\author{Saiful. R. Mondal}
\address{Department of Mathematics\\
King Faisal University, Al Ahsa 31982, Saudi Arabia}
\email{smondal@kfu.edu.sa}

\subjclass[2010]{26A33,33E20}
\keywords{Struve function, $\mathtt{k}$-Struve functions, fractional calculus, Pathway integral.}

\begin{abstract}
A new generalization called $\mathtt{k}$-Struve function and its properties given by Nisar and saiful very recently. In this paper, we establish the pathway fractional integral representation of  $\mathtt{k}$-Struve function. Many special cases also established to obtain the pathway integral representation of classical Struve function.  
\end{abstract}
\maketitle

\section{Introduction}
\label{Intro}

Recently, Nisar et al.\cite{Nisar-Saiful} introduced and studied various properties of $\mathtt{k}$-Struve function $\mathtt{S}_{\nu,c}^{\mathtt{k}}$ defined by

\begin{equation}\label{k-Struve}
\mathtt{S}_{\nu,c}^{\mathtt{k}}(x):=\sum_{r=0}^{\infty}\frac{(-c)^r}
{\Gamma_{\mathtt{k}}(r\mathtt{k}+\nu+\frac{3\mathtt{k}}{2})\Gamma(r+\frac{3}{2})}
\left(\frac{x}{2}\right)^{2r+\frac{\nu}{\mathtt{k}}+1}.
\end{equation}
where $c,\nu \in \mathbb{C}, \nu>\frac{3}{2}\mathtt{k}$.
The generalized Wright hypergeometric function ${}_{p}\psi _{q}(z)$ is given by
the series 
\begin{equation}
{}_{p}\psi _{q}(z)={}_{p}\psi _{q}\left[ 
\begin{array}{c}
(a_{i},\alpha _{i})_{1,p} \\ 
(b_{j},\beta _{j})_{1,q}%
\end{array}%
\bigg|z\right] =\displaystyle\sum_{k=0}^{\infty }\dfrac{\prod_{i=1}^{p}%
\Gamma (a_{i}+\alpha _{i}k)}{\prod_{j=1}^{q}\Gamma (b_{j}+\beta _{j}k)}%
\dfrac{z^{k}}{k!},  \label{Fox-Wright}
\end{equation}%
where $a_{i},b_{j}\in \mathbb{C}$, and real $\alpha _{i},\beta _{j}\in 
\mathbb{R}$ ($i=1,2,\ldots ,p;j=1,2,\ldots ,q$). Asymptotic behavior of this
function for large values of argument of $z\in {\mathbb{C}}$ were studied in 
\cite{CFox} and under the condition 
\begin{equation}
\displaystyle\sum_{j=1}^{q}\beta _{j}-\displaystyle\sum_{i=1}^{p}\alpha
_{i}>-1  \label{eqn-5-Struve}
\end{equation}%
was found in the work of \cite{Wright-2,Wright-3}. Properties of this
generalized Wright function were investigated in \cite{Kilbas}, (see also 
\cite{Kilbas-itsf, Kilbas-frac}. In particular, it was proved \cite{Kilbas}
that ${}_{p}\psi _{q}(z)$, $z\in {\mathbb{C}}$ is an entire function under
the condition ($\ref{eqn-5-Struve}$).

In\cite{Nair-1}, Nair introduced a pathway fractional integral operator and developed further by
Mathai and Haubold \cite{Mathai-Habold-1} ,\cite{Mathai-Habold-2} (Also, see \cite{Mathai-pathway}) is defined as follows :

Let $\ f\left( x\right) \in L\left( a,b\right) ,\eta \in \mathbb{C},\Re\left( \eta
\right) >0,a>0$ and the pathway parameter $\alpha <1$(cf \cite%
{Praveen-pathway}),then

\begin{equation}
\left( P_{0+}^{\left( \eta ,\alpha \right) }f\right) \left( x\right)
=x^{\eta }\int\limits_{0}^{\left[ \frac{x}{a\left( 1-\alpha \right)}\right]
}\left[1- \frac{a\left( 1-\alpha \right) t}{x}\right] ^{\frac{\eta }{\left(
1-\alpha \right)}}f\left( t\right) dt.  \label{eqn-path-1}
\end{equation}

For a real scalar $\alpha$, the pathway model for scalar random variables is
represented by the following probability density function (p.d.f.):

\begin{equation}
f\left( x\right) =c\left\vert x\right\vert ^{\gamma -1}\left[ 1-a\left(
1-\alpha \right) \left\vert x\right\vert ^{\delta }\right] ^{\frac{\beta }{%
\left( 1-\alpha \right) }},  \label{eqn-path-2}
\end{equation}

provided that $-\infty <x<\infty ,\delta >0,\beta \geq 0,\left[ 1-a\left(
1-\alpha \right) \left\vert x\right\vert ^{\delta }\right] >0,$ and $\gamma
>0$, where is the normalizing constant and $\alpha$ is called the pathway parameter \cite{Nair-1}.
 
%\begin{eqnarray*}
%c &=&\frac{1}{2}\frac{\delta \left[ a\left( 1-\alpha \right) \right] ^{\frac{%
%\gamma }{\delta }}\Gamma \left( \frac{\gamma }{\delta }+\frac{\beta }{%
%1-\alpha }+1\right) }{\Gamma \left( \frac{\gamma }{\delta }\right) \Gamma
%\left( \frac{\beta }{1-\alpha }+1\right) },\text{for }\alpha <1 \\
%&=&\frac{1}{2}\frac{\delta \left[ a\left( 1-\alpha \right) \right] ^{\frac{%
%\gamma }{\delta }}\Gamma \left( \frac{\beta }{\alpha -1}\right) }{\Gamma
%\left( \frac{\gamma }{\delta }\right) \Gamma \left( \frac{\beta }{\alpha -1}-%
%\frac{\gamma }{\delta }\right) },\text{for }\frac{1}{1-\alpha }-\frac{\gamma 
%}{\delta }>0,\alpha >1 \\
%&=&\frac{1}{2}\frac{\left( a\beta \right) ^{\frac{\gamma }{\delta }}}{\Gamma
%\left( \frac{\gamma }{\delta }\right) },\alpha \rightarrow 1
%\end{eqnarray*}

Note that for $\alpha <1$ it is a finite range density with $\left[
1-a\left( 1-\alpha \right) \left\vert x\right\vert ^{\delta }\right] >0$ and
\ ($\ref{eqn-path-2}$) remains in the extended generalized type-1 beta
family \ . The pathway density in ($\ref{eqn-path-2}$), for $\alpha < 1$,
includes the extended type-1 beta density, the triangular density, the
uniform density and many other p.d.f'.s .\cite{Praveen-pathway}.For instance
, $\alpha >1$ gives
\begin{equation}
f\left( x\right) =c\left\vert x\right\vert ^{\gamma -1}\left[ 1+a\left(
1-\alpha \right) \left\vert x\right\vert ^{\delta }\right] ^{-\frac{\beta }{%
\left( 1-\alpha \right) }},  \label{eqn-path-3}
\end{equation}%

provided that $-\infty <x<\infty ,\delta >0,\beta \geq 0,$ and $\alpha >0$
which is the extended generalized type-2 beta model for real x. It includes
the type-2 beta density, the F density, the Student-t density, the Cauchy
density and many more. For more details about pathway integral operator, one can refer \cite{Praveen-pathway, Purohit}.The purpose of this work is to investigate the composition formula of integral transform operator due to Nair, which is expressed in terms of the generalized Wright hypergeometric function, by inserting the $\mathtt{k}-$ Struve function 

\section{Pathway Fractional Integration of$\mathtt{k}$-Struve function.}

The results given in this section are based on the preliminary assertions
giving by composition formula of pathway fractional integral ($\ref%
{eqn-path-1}$) with a power function.

\begin{lemma}
({Agarwal}~\cite{Praveen-pathway},Lemma 1)

Let $\eta \in \mathbb{C},\Re\left( \eta \right) >0,\beta \in \mathbb{C}$ and $\alpha <1.$ If $%
\Re\left( \beta \right) >0,$and $\Re\left( \frac{\eta }{1-\alpha }\right) >-1,$%
then

\begin{equation}
\left\{ P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\beta -1}\right]
\right\} \left( x\right) =\frac{x^{\eta +\beta }}{\left[ a\left( 1-\alpha
\right) \right] ^{\beta }}\frac{\Gamma \left( \beta \right) \Gamma \left( 1+%
\frac{\eta }{1-\alpha }\right) }{\Gamma \left( 1+\frac{\eta }{1-\alpha }+\beta \right) }.
\label{lemma1}
\end{equation}
The pathway fractional integration of the $\mathtt{k}-$ Struve function is given by the following theorem.
\end{lemma}

\begin{theorem}\label{Th1}
Let $\eta ,\rho ,\nu, c \in C$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}\mathtt{k}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%

\begin{equation}\label{eqn1-th1}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\mathtt{S}_{\nu,c}^{\mathtt{k}}(t)\right]
\left( x\right) =x^{\eta}\left(\frac{x}{a(1-\alpha)}\right)^{\rho+\frac{\nu}{\mathtt{k}}+1}\frac{\Gamma\left(1+\frac{\eta}{1-\alpha}\right)}
{\mathtt{k}^{\frac{\nu}{\mathtt{k}}+\frac{1}{2}}2^{\frac{\nu}{\mathtt{k}}+1}}\\ 
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +\frac{\nu}{\mathtt{k}}+1,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +\frac{\nu}{\mathtt{k}}+\frac{\eta }{1-\alpha }+2,2\right) , & \left(\frac{\nu}{\mathtt{k}}+\frac{3}{2},1\right) , & \left( 3/2,1\right)%
\end{array}%
;-\frac{cx^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right].%
\end{array}
\end{equation}
\end{theorem}

\begin{proof}
Applying the pathway operator defined in \eqref{eqn-path-1} to \eqref{k-Struve}, and changing the order of integration and summation, we get%

\begin{align*}
\left( P_{0+}^{\left( \eta ,\alpha \right)}\left[ t^{\rho-1}\mathtt{S}_{\nu,c}^{\mathtt{k}}(t)\right] \right)
\left( x\right)&=P_{0+}^{\left( \eta ,\alpha \right)}\left[t^{\rho-1}\sum_{r=0}^{\infty}\frac{(-c)^{r}\left(\frac{t}{2}\right)^{2r+\frac{\nu}{\mathtt{k}}+1}}{\Gamma_{\mathtt{k}}\left(r\mathtt{k}+\nu+\frac{3}{2}\mathtt{k}\right)\Gamma\left(r+\frac{3}{2}\right)}\right](x)\\
&=\sum_{r=0}^{\infty}\frac{(-c)^{r}\left(\frac{1}{2}\right)^{2r+\frac{\nu}{\mathtt{k}}+1}}
{\Gamma_{\mathtt{k}}\left(r\mathtt{k}+\nu+\frac{3}{2}\mathtt{k}\right)\Gamma\left(r+\frac{3}{2}\right)}
P_{0+}^{\left( \eta ,\alpha \right)}\left(t^{\rho+2r+\frac{\nu}{\mathtt{k}}}\right)(x)
\end{align*}
Using Lemma $(\ref{lemma1})$, we get
\begin{align*}
&&=\sum_{r=0}^{\infty}\frac{(-c)^{r}\left(\frac{1}{2}\right)^{2r+\frac{\nu}{\mathtt{k}}+1}}
{\Gamma_{\mathtt{k}}\left(r\mathtt{k}+\nu+\frac{3}{2}\mathtt{k}\right)\Gamma\left(r+\frac{3}{2}\right)}
\frac{x^{\eta+\rho+2r+\frac{\nu}{\mathtt{k}}+1}}{[a(1-\alpha)]^{\rho+2r+\frac{\nu}{\mathtt{k}}+1}}\\
&&\times\frac{\Gamma\left(\rho+2r+\frac{\nu}{\mathtt{k}}+1\right)\Gamma\left(1+\frac{\eta}{1-\alpha}\right)}{\Gamma\left(\frac{\eta}{1-\alpha}+\rho+2r+\frac{\nu}{\mathtt{k}}+2\right)}
\end{align*}
Now using the relation $\Gamma_{\mathtt{k}}\left(\gamma\right)=\mathtt{k}^{\frac{\gamma}{\mathtt{k}}-1}\Gamma\left(\frac{\gamma}{\mathtt{k}}\right)$, we get
\begin{align*}
&=\frac{x^{\eta+\rho+\frac{\nu}{\mathtt{k}}+1}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)}{\left[a(1-\alpha)\right]^{\rho+\frac{\nu}{\mathtt{k}}+1}2^{\frac{\nu}{\mathtt{k}}+1}}\\
&\times \sum_{r=0}^{\infty}\frac{(-c)^{r}x^{2r}}
{\mathtt{k}^{r+\frac{\nu}{\mathtt{k}}+\frac{1}{2}}\Gamma\left(r+\frac{\nu}{\mathtt{k}}+\frac{3}{2}\right)\Gamma\left(r+\frac{3}{2}\right)4^{r}[a(1-\alpha)]^{2r}}\\
&\times\frac{\Gamma\left(\rho+\frac{\nu}{\mathtt{k}}+1+2r\right)}{\Gamma(\frac{\eta}{1-\alpha}+\rho+2r+\frac{\nu}{\mathtt{k}}+2)}.
\end{align*}
In view of $(\ref{Fox-Wright})$, we arrived the desired result.
\end{proof}
\begin{corollary}
If we take $\mathtt{k}=1$ in theorem \eqref{Th1}, then we get the pathway integrals involving classical Struve function as:
\begin{equation}\label{eqn1-cor1}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\mathtt{S}_{\nu,c}^{1}(t)\right]
\left( x\right) =x^{\eta}\left(\frac{x}{a(1-\alpha)}\right)^{\rho+\nu+1}
\frac{\Gamma\left(1+\frac{\eta}{1-\alpha}\right)}{2^{\nu+1}}\\ 
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +\nu+1,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +\nu+\frac{\eta }{1-\alpha }+2,2\right) , & \left(\nu+\frac{3}{2},1\right) , & \left( 3/2,1\right)%
\end{array}%
;-\frac{cx^{2}}{4\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right].%
\end{array}
\end{equation}
\end{corollary}

Now, we will give the relation relation between trigonometric function and $\mathtt{k}$-Struve function. By taking $\nu=\mathtt{k}/2$ 
in (3.10) of \cite{Nisar-Saiful} we get the relation between cosine functions and $\mathtt{k}$-Struve functions as
\begin{equation}\label{cos}
1-\cos\left(\frac{\alpha x}{\sqrt{\mathtt{k}}}\right)= \frac{\alpha}{\mathtt{k}}\sqrt{\frac{\pi x}{2}}~\mathtt{S}_{\frac{\mathtt{k}}{2}, \alpha^2}^{\mathtt{k}} (x).
\end{equation}

Similarly, the relation

\begin{equation}\label{cosh}
\cosh\left(\frac{\alpha x}{\sqrt{\mathtt{k}}}\right)-1= \frac{\alpha}{\mathtt{k}}\sqrt{\frac{\pi x}{2}}~\mathtt{S}_{\frac{\mathtt{k}}{2}, -\alpha^2}^{\mathtt{k}} (x),
\end{equation}
can be derive from (3.11) of \cite{Nisar-Saiful}.

Also, by taking $\nu=-\frac{k}{2}$ in $(\ref{k-Struve})$, we obtained the following:
\begin{equation}\label{sin}
\sin\left(\frac{\alpha x}{\sqrt{\mathtt{k}}}\right)=\alpha\left(\sqrt{\frac{\pi x}{2\mathtt{k}}}\right)\mathtt{S}_{-\frac{\mathtt{k}}{2},\alpha^{2}}^{\mathtt{k}}\left(x\right).
\end{equation}

\begin{equation}\label{sinh}
\sinh\left(\frac{\alpha x}{\sqrt{\mathtt{k}}}\right)=\alpha\left(\sqrt{\frac{\pi x}{2\mathtt{k}}}\right)\mathtt{S}_{-\frac{\mathtt{k}}{2},-\alpha^{2}}^{\mathtt{k}}\left(x\right).
\end{equation}

\section{Pathway fractional integration of cosine,hyperbolic cosine, sine
and hyperbolic sine functions}
\begin{theorem}\label{Th2}
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}\mathtt{k}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%

\begin{equation}\label{eqn1-th2}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(1-\cos\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
\left( x\right) =\sqrt{\pi}\frac{\gamma}{\mathtt{k}^{2}}\frac{ x^{\eta+\rho+2}}{4[a(1-\alpha)]^{\rho+2}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +2,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +3+\frac{\eta }{1-\alpha },2\right) , & \left(2,1\right) , & \left( 3/2,1\right)%
\end{array}%
;-\frac{\gamma^{2}x^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{theorem}

\begin{proof}
Applying the pathway operator defined in \eqref{eqn-path-1} to \eqref{cos}, and changing the order of integration and summation, we get%
\begin{align*}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(1-\cos\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
&=\left( P_{0+}^{\left( \eta ,\alpha \right)}\left[ t^{\rho-1}\frac{\gamma}{\mathtt{k}}\sqrt{\frac{\pi t}{2}}\mathtt{S}_{\frac{k}{2},\gamma^{2}}^{\mathtt{k}}(t)\right] \right)\left( x\right)\\
&=\sqrt{\frac{\pi}{2}}\frac{\gamma}{\mathtt{k}}\sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}\left(\frac{1}{2}\right)^{2r+\frac{1}{2}+1}}{\Gamma_{\mathtt{k}}\left(r\mathtt{k}+\frac{4}{2}\mathtt{k}\right)\Gamma\left(r+\frac{3}{2}\right)}P_{0+}^{\left( \eta ,\alpha \right)}\left[t^{\rho+2r+1}\right](x),
\end{align*}
Using Lemma \ref{lemma1}, we get
\begin{align*}
=\sqrt{\pi}\frac{\gamma}{\mathtt{k}}\sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}(\frac{1}{2})^{2r+2}}{\Gamma_{\mathtt{k}}(r\mathtt{k}+2\mathtt{k})\Gamma(r+\frac{3}{2})}\frac{x^{\eta+\rho+2+2r}}{[a(1-\alpha)]^{\rho+2+2r}}\\
\times \frac{\Gamma(\rho+2r+2)\Gamma(1+\frac{\eta}{1-\alpha})}{\Gamma(1+\frac{\eta}{1-\alpha}+\rho+2+2r)}
\end{align*}
Now using the relation $\Gamma_{\mathtt{k}}\left(\gamma\right)=\mathtt{k}^{\frac{\gamma}{\mathtt{k}}-1}\Gamma\left(\frac{\gamma}{\mathtt{k}}\right)$, we get
\begin{align*}
&=\sqrt{\pi}\frac{\gamma}{\mathtt{k}^{2}}\frac{ x^{\eta+\rho+2}}{4[a(1-\alpha)]^{\rho+2}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
&=\sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}(\frac{1}{2})^{2r}\Gamma\left(\rho+2+2r\right)}{\Gamma(r+2)\Gamma(r+\frac{3}{2})}\frac{x^{2r}}{[a(1-\alpha)]^{2r}\Gamma(1+\frac{\eta}{1-\alpha}+\rho+2+2r)}
\end{align*}
In view of $(\ref{Fox-Wright})$, we arrived the desired result.
\end{proof}

\begin{corollary}\label{Cor2}
If we take $\mathtt{k}=1$ in theorem \ref{Th2}, then we get the pathway integrals involving classical Struve function as:
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold%
\begin{equation}\label{eqn1-Cor2}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right)}\left[t^{\rho-1}\left(1-cos{\gamma t}\right)\right]\left( x\right) =\sqrt{\pi}{\gamma}\frac{x^{\eta+\rho+2}}{4[a(1-\alpha)]^{\rho+2}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +2,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +3+\frac{\eta }{1-\alpha },2\right) , & \left(2,1\right) , & \left( 3/2,1\right)%
\end{array}%
;-\frac{\gamma^{2}x^{2}}{4\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{corollary}

\begin{theorem}\label{Th3}
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}\mathtt{k}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-th3}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\cosh\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
\left( x\right) =\sqrt{\pi}\frac{\gamma}{\mathtt{k}^{2}}\frac{ x^{\eta+\rho+2}}{4[a(1-\alpha)]^{\rho+2}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +2,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +3+\frac{\eta }{1-\alpha },2\right) , & \left(2,1\right) , & \left( 3/2,1\right)%
\end{array}%
;\frac{\gamma^{2}x^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{theorem}

\begin{corollary}\label{cor3}
If we set $\mathtt{k}=1$ in Theorem \ref{Th3} then we get,
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-Cor3}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\cosh\left(\gamma t\right)\right)\right]
\left( x\right) =\sqrt{\pi}\gamma\frac{ x^{\eta+\rho+2}}{4[a(1-\alpha)]^{\rho+2}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{2}\Psi _{3}\left[ 
\begin{array}{ccc}
\left( \rho +2,2\right) , & \left( 1,1\right); &  \\ 
\left( \rho +3+\frac{\eta }{1-\alpha },2\right) , & \left(2,1\right) , & \left( 3/2,1\right)%
\end{array}%
;\frac{\gamma^{2}x^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{corollary}

\begin{theorem}\label{Th4}
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}\mathtt{k}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-th4}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\sin\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
\left( x\right) =\gamma\sqrt{\frac{\pi}{\mathtt{k}}}\frac{x^{\rho+\eta+1}}{2}{\left[a(1-\alpha)\right]^{\rho+\frac{1}{2}}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{1}\Psi _{2}\left[ 
\begin{array}{ccc}
\left( \rho +\frac{1}{2},2\right) ; &  \\ 
\left( \rho +\frac{\eta }{1-\alpha }+\frac{3}{2},2\right) , & \left(\frac{3}{2},1\right); & %
\end{array}%
;\frac{-\gamma^{2}x^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{theorem}

\begin{proof}
Applying the pathway operator defined in \eqref{eqn-path-1} to \eqref{cos}, and changing the order of integration and summation, we get%
\begin{align*}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\sin\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
&=\left( P_{0+}^{\left( \eta ,\alpha \right)}\left[ t^{\rho-1}\gamma\sqrt{\frac{\pi x}{2\mathtt{k}}}~\mathtt{S}_{-\frac{\mathtt{k}}{2},\gamma^{2}}^{\mathtt{k}}(t)\right] \right)\left( x\right)\\
&=\gamma\sqrt{\frac{\pi x}{2\mathtt{k}}}P_{0+}^{\left( \eta ,\alpha \right)}\left[t^{\rho-1}\sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}
\left(\frac{t}{2}\right)^{2r-\frac{1}{2}+1}}{\Gamma_{\mathtt{k}}(r\mathtt{k}-\frac{\mathtt{k}}{2}+\frac{3\mathtt{k}}{2})\Gamma\left(r+\frac{3}{2}\right)}\right](x)\\
&=\gamma\sqrt{\frac{\pi x}{2\mathtt{k}}}\sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}\left(\frac{1}{2}\right)^{2r+\frac{1}{2}}}{\Gamma_{\mathtt{k}}\left(r\mathtt{k}+\mathtt{k}\right)\Gamma\left(r+\frac{3}{2}\right)}P_{0+}^{\left( \eta ,\alpha \right)}\left[t^{\rho+2r+\frac{1}{2}-1}\right]
\end{align*}
Using Lemma \ref{lemma1} and the relation $\Gamma_{\mathtt{k}}\left(\gamma\right)=\mathtt{k}^{\frac{\gamma}{\mathtt{k}}-1}\Gamma\left(\frac{\gamma}{\mathtt{k}}\right)$, we get
\begin{align*}
&=\gamma\sqrt{\frac{\pi}{\mathtt{k}}}\frac{x^{\rho+\eta+1}}{[a(1-\alpha)]^{\rho+\frac{1}{2}}}\\
&\times \sum_{r=0}^{\infty}\frac{(-\gamma^{2})^{r}\left(\frac{1}{2}\right)^{2r}x^{2r}\Gamma(\rho+\frac{1}{2}+2r)}{\Gamma\left(r+1\right)\Gamma\left(r+\frac{3}{2}\right)\Gamma\left(\rho+\frac{\eta}{1-\alpha}+\frac{1}{2}+1+2r\right)k^{r}[a(1-\alpha)]^{2r}}
\end{align*}
In view of $(\ref{Fox-Wright})$, we arrived the desired result.
\end{proof}

\begin{corollary}\label{Cor4}
If we take $\mathtt{k}=1$, then we have
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-Cor4}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\sin\left(\gamma t\right)\right)\right]
\left( x\right) =\gamma\sqrt{\pi}\frac{x^{\rho+\eta+1}}{2}{\left[a(1-\alpha)\right]^{\rho+\frac{1}{2}}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{1}\Psi _{2}\left[ 
\begin{array}{ccc}
\left( \rho +\frac{1}{2},2\right) ; &  \\ 
\left( \rho +\frac{\eta }{1-\alpha }+\frac{3}{2},2\right) , & \left(\frac{3}{2},1\right); & %
\end{array}%
;\frac{-\gamma^{2}x^{2}}{4\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{corollary}

\begin{theorem}\label{Th5}
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}\mathtt{k}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-th5}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\sinh\left(\frac{\gamma t}{\sqrt{\mathtt{k}}}\right)\right)\right]
\left( x\right) =\gamma\sqrt{\frac{\pi}{\mathtt{k}}}\frac{x^{\rho+\eta+1}}{2}{\left[a(1-\alpha)\right]^{\rho+\frac{1}{2}}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{1}\Psi _{2}\left[ 
\begin{array}{ccc}
\left( \rho +\frac{1}{2},2\right) ; &  \\ 
\left( \rho +\frac{\eta }{1-\alpha }+\frac{3}{2},2\right) , & \left(\frac{3}{2},1\right); & %
\end{array}%
;\frac{\gamma^{2}x^{2}}{4\mathtt{k}\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{theorem}

\begin{corollary}\label{Cor5}
If we take $\mathtt{k}=1$, then we have
Let $\eta ,\rho ,\nu, c \in \mathbb{C}$ and $\alpha <1$ be such that $\Re\left(
\eta \right) >0,\Re\left( \rho \right) >0, \nu>-\frac{3}{2}$ and $\Re\left( \frac{\eta }{1-\alpha 
}\right) >-1 $ then the following formula hold true%
\begin{equation}\label{eqn1-Cor5}
\begin{array}{c}
P_{0+}^{\left( \eta ,\alpha \right) }\left[ t^{\rho-1}\left(\sinh\left(\gamma t\right)\right)\right]
\left( x\right) =\gamma\sqrt{\pi}\frac{x^{\rho+\eta+1}}{2}{\left[a(1-\alpha)\right]^{\rho+\frac{1}{2}}}\Gamma\left(1+\frac{\eta}{1-\alpha}\right)\\
\times _{1}\Psi _{2}\left[ 
\begin{array}{ccc}
\left( \rho +\frac{1}{2},2\right) ; &  \\ 
\left( \rho +\frac{\eta }{1-\alpha }+\frac{3}{2},2\right) , & \left(\frac{3}{2},1\right); & %
\end{array}%
;\frac{\gamma^{2}x^{2}}{4\left[ a^{2}\left( 1-\alpha \right)^{2} \right]}\right]%
\end{array}
\end{equation}
\end{corollary}

\end{document}